# Unsocial Intelligence: a Pluralistic, Democratic, and Participatory Investigation of AGI Discourse

### v2

@@ 5. Conclusion: democratic intelligence
Building on John Dewey’s conception of “democracy as the use of social intelligence to solve problems of practical interest”, Anderson (2006) argues that democratic institutions embody solving social problems without resorting to oppression or coercion. Anderson argues that democracies are uniquely suited to investigating solutions to public interest problems, problems that need to be solved through deliberation (“votes and talk”) rather than through procedures like markets. They do so by pairing sources of legitimacy like procedural fairness and universal inclusion (through mechanisms like law, rights, and voting), experimentation (“revising […] decisions on the basis of experience with their consequence”), and dissent (Anderson, 2006).
@@
Putnam (2011) summarizes this view: democracy “is the precondition for the full application of intelligence to the solution of social problems”. [^15] See Festenstein (2019) for an overview of critiques of this family of views, both as interpretations of Dewey and as conceptions of democracy. This lens puts pressure on the narrow range of problems, tasks, and processes that take center stage in discussions of would-be human-level AI and AGI. Intelligent solutions to social problems should not be framed as a matter of finding optimal means for satisfying fixed preferences. Rather, they require procedures that support universal inclusion in interrogation, deliberation, and dissent about what counts as the “common good” (Putnam, 2011) and the “public interest” (Anderson, 2006). They require procedures with political legitimacy (Peter, 2017). On this view, would-be human-level AI or AGI can provide intelligent solutions to social problems only if they are also objects of dissent and deliberative contestation, rather than systems designed from the top-down by “experts” only. The project of imagining “worlds” (Costanza-Chock, 2020) and socially transformative technologies worth building should be one of collective “experiments in living” (Mill, 2003; Anderson, 1991), in which all impacted rights-holders hold decision-makers accountable.
@@ Acknowledgements.
6. Positionality and ethics statement
The authors of this work take a humanistic lens—drawing especially on feminist philosophy of science, social philosophy, and STS—to a deeply cross-disciplinary topic. In particular, we see ourselves as having an outsider perspective on both human intelligence research (such as psychometrics) and AGI discourse. We see bringing a critical lens to practices that are not our own as running the risk of becoming an assertion of power: imposing our own experiences and beliefs upon the practices and beliefs of others. We can neither claim to share, nor to fully appreciate the perspectives of the many communities whose beliefs we are striving to interrogate.
We are also practitioners with industry and civil society ties to responsible AI development. As practitioners, we see the growing importance of the topic of AGI and human-level AI in private sector, civil society, and government discussions of AI governance as requiring interrogation from a broader range of perspectives. In particular, we are concerned with what we perceive as the narrow range of communities that currently play a dominant role in shaping discourse on this topic.
We take seriously the idea that the topic of human-level AI or AGI concerns potentially socially transformative technology. We take a democratic, contextualist, sociotechnical, and participatory perspective on the topic of transformative technologies: as requiring processes that empower inclusive deliberation and dissent going far beyond the perspectives of the handful of expert communities who currently play a predominant role in shaping their development. Socially transformative technologies affect everyone. We see this project as highlighting decision points and assumptions that are important to interrogate in bringing democratic and participatory deliberation to bear on the topic of transformative forms of machine intelligence.
@@ Appendix A Dimensions of AGI: a summary
[ label = none, entry = none, ] width = colspec = Q[100]$Q[347]$, hlines, vlines Dimension & Description Examples
Embodiment Whether the ability to accomplish certain physical tasks is part of the definition Required: (Wozniak, 2010; Marcus, 2022; Weizenbaum, 1976)

### v3

@@ 5. Conclusion: politically legitimate intelligence
Building on John Dewey’s conception of “democracy as the use of social intelligence to solve problems of practical interest”, Anderson (2006) argues that democratic institutions embody solving social problems without resorting to oppression or coercion. She argues that democracies are uniquely suited to investigating solutions to public interest problems, problems that need to be solved through deliberation (“votes and talk”) rather than through procedures like markets. They do so by pairing sources of legitimacy like procedural fairness and universal inclusion (through mechanisms like law, rights, and voting), experimentation (“revising […] decisions on the basis of experience with their consequence”), and dissent (Anderson, 2006).
@@
Putnam (2011) summarizes this view: democracy “is the precondition for the full application of intelligence to the solution of social problems”. [^15] See Festenstein (2019) for an overview of critiques of this family of views, both as interpretations of Dewey and as conceptions of democracy. This lens puts pressure on the narrow range of problems, tasks, and processes that take center stage in discussions of would-be human-level AI and AGI. Intelligent solutions to social problems should not be framed as a matter of finding optimal means for satisfying fixed preferences. Rather, they require procedures that support universal inclusion in interrogation, deliberation, and dissent about what counts as the “common good” (Putnam, 2011) and the “public interest” (Anderson, 2006). They require procedures with political legitimacy (Peter, 2017). On this view, future forms of AI can provide intelligent solutions to social problems only if they are also objects of dissent and deliberative contestation, rather than systems designed from the top-down by “experts” only. The project of imagining “worlds” (Costanza-Chock, 2020) and future technologies worth building should be one of collective “experiments in living” (Mill, 2003; Anderson, 1991), in which all impacted rights-holders hold decision-makers accountable.
@@ Appendix A Dimensions of AGI: A Summary
[ caption = Summary table of how different conceptions of AGI differ on value-laden dimensions, label = dimensions, entry = none, ] width = colspec = Q[100]Q[347]Q[347], hlines, vlines Dimension & Description Examples
Embodiment Whether the ability to accomplish certain physical tasks is part of the definition Required: (Fast Company, 2010; Marcus, 2022; Weizenbaum, 1976)